X + 2y - Z = 0
3x + y -2z = 3
2x - z = 0

Answer by Alan3354(69443)

X + 2y - Z = 0
3x + y -2z = 3
2x - z = 0 --> z = 2x
--------------
Sub for z in eqn 1 & 2
---
X + 2y - Z = 0 --> -x + 2y = 0
3x + y -2z = 3 --> -x + y = 3
------
-x + 2y = 0
-x + y = 3
-------------------- Subtract
y = -3
=====================
-x + y = 3
x = -6
=====================
z = -12